Given the following system of equations, identify the type of system.
x + y = 4
x - y = 6

Answer by ewatrrr(24785)

 
Hi
A consistent system has at least one solution.
An inconsistent system has no solutions.
An independent system has exactly one solution.
A dependent system has infinitely many solutions.
x + y = 4
x - y = 6 | Adding 1st and 2nd EQs to eliminate y
2x = 10
x = 5 and y = -1 | 5 + y = 4, y = - 1
Independent System (one solution)
CHECKING our answer***
Using x - y = 6
5 - (-1) = 6
Wish You the Best in your Studies.